OpenStudy (anonymous):

Calculate the wavelength, λ, in meters of a photon capable of exciting an electron in He+ from the ground state to n = 4.

OpenStudy (anonymous):

\[\frac{ 1 }{ \lambda }= R(\frac{ 1 }{ S^{2} } - \frac{ 1 }{ L ^{2} })\] Where R is Rydbergs constant S is the level original quantum number, L is the quantum number changed to